Subject: 3.5-m Users Committee minutes 4/16/07

From: Bruce Gillespie

Submitted: Thu, 19 Apr 2007 10:46:47 -0600

Message number: 1027 (previous: 1026, next: 1028 up: Index)

                         APO 3.5-m Users Committee Phonecon, 4/16/07

Attending:  Suzanne Hawley, Bruce Gillespie, John Wilson (for Remy  
Indebetouw), Russet McMillan, Jon Fulbright, Mark Klaene, John Bally,  
Jon Holtzman, Scott Anderson

Absent: Michael Strauss, Al Harper

Minutes taken by Bruce Gillespie


Bruce began with a side-bar announcement:  Since the Futures  
Committee report is now finished, Don York has asked to  absent  
himself from the Users Committee.  We thank Don for all of his time  
and involvement with the Users Committee over the past years.


User feedback, comments from institutional representatives:

Princeton - (Strauss) [e-mailed from Michael: "I'm happy to report  
that we've had no problems with observing over the last month."]

JHU (Fulbright) - Jon said he had no user comments.

Chicago (Harper) - no report.

Washington (Anderson) - Scott said he had a spectacular run this past  
Saturday, with better than .6 arcsecond seeing in I band.  Suzanne  
added that she recently was observing at APO in person (using Agile),  
while a student at UW was helping remotely, all with support from  
three Observing Specialists (who happened to all be at APO for other  
purposes).  The run was quite successful thanks to the excellent  

NMSU - (Holtzman) - Jon said that the recent Pluto occultation  
observations were weathered out.

Virginia (Wilson) - John said that Anne Verbiscer recently had her  
first remote observing run, and she wanted to thank the Observing  
Specialists for helping to make it successful.

Colorado (Bally) - John had nothing to report from the users.  He did  
announce that the two NIC-FPS detectors (engineering- and science- 
grade) were going to be sent to Rockwell to rebond the leads to  
enable 16-channel reads.  Rockwell has agreed to do this at no cost.   
Also, an order has been placed with IR Labs for the Fabry-Perot cryo  
test dewar.  Also, a new Leach controller is being ordered for the  
testing of the F-P this summer; in addition, this controller will  
serve as a backup for NIC-FPS and TripleSpec.


Discussion of telescope/instruments report - Klaene:

The report is given below; there was some discussion:

Mark added that yesterday the DIS blue camera started warming up.   
Pumping on the camera and tapping on the ion pump seemed to fix the  
problem.  Russet added that there didn't appear to be any new  
residual scattered light in the Blue images taken last night.

On the drive replacement item, Russet said that in recent weeks the  
tracking in the western and southern parts of the sky has been a  
little ratty, usually only seen in NIC-FPS exposures since it  
typically yields the best image quality.  She has adjusted the  
tracking parameters, and Mark has cleaned the az drives recently, so  
if any users see signs of tracking errors they should immediately let  
their Observing Specialist know.  Suzanne asked if all three boxes  
are to be swapped in May, and Mark said that is the plan.  Bruce  
added that we are looking into collecting image quality data, and  
compare it with before and after the installation of the new top  
end.  Russet noted that the collimation of the telescope now seems to  
be very solid, compared to how it was before the new top end was  


   3.5-m Telescope, Instruments, and CIF Projects Highlights, 3/17/07  
through 4/13/07

                                Mark Klaene

0) Overview

Spring weather intermixed with mild winter storms have been the story  
this past month.  Winds are increasing along with the background dust  
levels. The week of April 2nd included numerous meetings on site with  
design reviews for the rotator upgrade and the new axis controllers  
along with software reviews and an operations staff meeting.

1) Telescope

No significant problems this period. The engineering time in April  
was used to upgrade the 2 transverse actuator motors on the secondary  
and replace the B actuator assembly. The B actuator load limit  
switches were adjusted and appeared to correct the error were seeing  
during focus adjustments.

All 3 replacement spare drive boxes have been assembled and are ready  
to be swapped in during the May engineering time.

2) Instruments

Agile completed it's proof of concept run; the camera used was a  
loaner and not the actual camera purchased.  This was due to a failed  
thermoelectric cooler in the original camera. That camera has been  
repaired and is on its way to UW for testing.  Various hardware  
modifications to the mounting system are also in work as well as  
testing with the fiber-optic interface.

A replacement flip mirror motor was obtained for the Echelle.  
Replacement will have to wait until next month as the replacement  
motor is not a direct replacement and will require re-wiring.

APOLLO interlock issues are continuing and being worked on by the  
APOLLO group.

A modified (recommended via a study by JHU) grating holder for the  
DIS gratings is being tested with the hope that the modification will  
ease installation and lower forces on the glass.

A decision to purchase a spare Leach controller was made for NIC-FPS  
and TripleSpec. The purchase at this point will allow CU to test the  
engineering grade chip in 16-channel readout mode without taking NIC- 
FPS out of service this summer.

3) Engineering and CIF projects

A decision to proceed with the tertiary support  project was made  
although engineering resources are limited until Jan 2008.  Most of  
the parts are on hand for phase one of the axis controllers. Much of  
the software is completed and running in a simulation mode. Plan is  
to integrate it with the rotator first this summer.  Servo motor  
selection is underway for the rotator and other axes.

4) Miscellaneous

Truss calibration lamp replacement project is still in the research  
phase. Work on purge air and heat exhaust is being done on site to  
prepare for the TripleSpec arrival. Preparations are continuing for  
3.5m re-aluminization this summer.  Summer shutdown schedules are  
firmed up, planned for Aug 20 - Sept 16.  TUI 1.3 was released.


Engineering Reviews - Klaene/Hawley:

Mark reported on the recent in-person engineering discussions/reviews  
held at APO.  Mike Skrutskie, John Wilson, Matt Nelson, Russell Owen,  
Mike Carr, Suzanne Hawley, and Larry Carey participated in these  
meetings with the APO day and night staff, which helped to make them  
very productive:

   o Software for TripleSpec--integration and operational issues were  
discussed. [TripleSpec is targeted for delivery at APO in October]

   o CDR for rotator upgrade design by Mike Carr--the review led to a  
few suggestions, some which will be implemented and others still in  
discussion.  A newer-generation motor concept is being considered.

   o New axis controllers design by Fritz Stauffer--the design  
appears to be quite well along, and trial runs are planned for this  

   o Direct drives concept study--in the hope of eventually replacing  
our drive boxes (to improve pointing/tracking performance and  
maintenance requirements), a study has started for conversion to a  
direct-drive system for the az and alt motions.  A simple approach of  
using a design and hardware proposed fifteen years ago has been found  
to be inadequate.  Newer design approaches are being studied, but  
final decisions are not expected to be taken until later this year.

   o Agile future--assuming commissioning is successful, Agile would  
move to a mirror-cell port, with a filter wheel and image rotator.   
Work on this will continue through the fall.

   o Tertiary upgrade--this is to replace the mounts and actuators on  
the tertiary mirror.  The engineering work has started this year,  
with a possible implementation next year.

Bruce added that the major upgrades planned for the telescope this  
year are the new axis controllers and the rotator upgrade--all the  
rest are in study and design phase.  Suzanne said that the major  
instrument work this year will be on Agile and Triplespec.  She also  
said that during the course of the reviews, she noted that there is a  
very fine group of systems engineering people at work on the various  
projects for the telescope and instruments, and that it was good that  
the Observing Specialists were participating in the design  
activities, lending operational perspectives where needed.  Once the  
summer shutdown is done, we will implement the rotator, and Agile  
upgrades and receive TripleSpec for commissioning.  If so decided,  
the direct drive upgrade will be started in 2008.


Future instruments for the 3.5-m - Hawley:

We're planning to integrate SPIcam into TUI this summer.  But it's  
apparent that we're eventually going to need a new visible imager,  
even though Agile (w/ focal reducer) and DIS imaging modes will  
somewhat give us a backup capability.  SPIcam is not well documented,  
and doesn't have a detector that performs to today's standards.   
Given that it takes several years to plan and build a new instrument,  
it is now the time to begin thinking of the next next-generation  
instrument for the 3.5-m, be it a visible imager or whatever.  Users  
Committee members should begin trolling their users on thoughts about  
a new instrument, which might be partially funded by CIF funds.   
Incidentally, we have recently seen proposals for external funding  
for upgrades to the Goddard Fabry-Perot and for high-precision radial- 
velocity instruments to be placed at the 3.5-m.


Actions from earlier meetings - group

[open from previous months]:

ACTION:  Rene to collect history of 3.5-m publications, and Bruce  
will make a space for them on the APO website.
STATUS:  Gretchen and Don York completed a draft of the 3.5-m  
publications list. It will be added as an appendix to the Futures  
Committee report, and also placed separately in a stand-alone web  
location (URLs will be posted).  Members of the Users Committee are  
invited to check the list against known publications from their  
respective institutions.

DISCUSSION:  The 3.5-m publications list is now posted at http://, and also at http://  Users are invited to  
look at the list to see if there are any residual errors, omission,  
or duplications.  This action is closed.

[new actions from last month]:

ACTION: We need some tests to determine the accuracy of the [DIS  
grating] moves so users could evaluate if not checking after the move  
in special situations would be acceptable.
STATUS: Mark Klaene reports "once calibrated, the repeatability of  
the motors during a move is to within 1 step. This should yield and  
overall accuracy of +/- 2.5 steps."  This should close this action,  
unless there is further discussion needed.

DISCUSSION: If you use the same grating and don't move the turret,  
this prescription seems to be correct.  But if gratings are changed  
or the turret switched, the wavelength calibration can be  
significantly different, and should be checked if the science  
requires it.  Jon Holtzman agreed to put some words to this effect in  
the DIS users documentation.  Suzanne asked about the status of the  
instrument on-line instrument documentation.  Jon said that the DIS  
documentation is pretty much done and on line, and the Echelle and  
NIC-FPS documentation is now being worked on.

ACTION: Suzanne will follow up with Russell Owen about TUI  
scripting.  Users are encouraged to provide feedback to their users  
reps about integrating Spicam into TUI.
STATUS: Russell is willing to include some modest number of useful  
TUI scripts when SPIcam is implemented into TUI.  No feedback has  
been received, so we plan to close this action item.

DISCUSSION:  Jon Holtzman will contact Russell to discuss the details  
of TUI scripts for SPIcam.  This action is closed.


On other topics, Suzanne mentioned that the summer shutdown is  
between 20 August and 16 September.  It is later than usual due to  
the timing of the aluminizing slot available at Kitt Peak that allows  
the SDSS telescope to get their primary coated earlier in the  
summer.  JonH said he would contact Josh Frieman about this schedule  
since it impacts the fall SDSS-II SNe follow-up observing with the  
3.5-m telescope.

Suzanne also mentioned that Jim Davenport, who is about to receive  
his B.S. from UW, will be working as an Assistant Observing  
Specialist on the 3.5-m this summer, helping to alleviate potential  
staffing shortfalls and to provide assistance with the shutdown.  Jim  
helped work on the installation of the new top end last summer.


Next meeting:

The next Users Committee phonecon will be on Monday, 14 May, at 8:30  
AM Pacific Time.  Agenda and other materials will be sent to the  
committee members during the preceding week.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1027 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to